# Non-oscillatory entropy stable DG schemes for hyperbolic conservation law

### v1

@@ 3.1 Integral factor SSP-RK method
The PDE can be solved in the MOL framework, that is after DG discretization the semi-discrete scheme can be treated as an ODE system, and then, we can apply a known ODE solver on it to obtain the fully discrete scheme. However, for the numerical solution containing discontinuities, the additional viscous term is a stiff term, leading a very strictly limited time step when coupling with explicit time discretization. To avoid this problem, we will employ the integrating factor strong-stability-preserving Runge-Kutta (IF-SSP-RK) method introduced in [6]. Here, we only discuss the one-dimensional scalar case for simplify.
@@
To introduce the IF-SSP-RK, we begin with a specific case that this stiff term is semilinear, i.e. $\mathcal{L}^{v}_{h}(\mathbf{u}_{h})=\mathbf{C}\mathbf{u}_{h}$, where $\mathbf{C}$ is a constant matrix. Then we can multiply the factor $e^{-\mathbf{C}t}$ on both sides and get that
@@ 3.2 Positivity-preserving limiter
For some extreme problems, the negative density or pressure may be introduced. If the density or pressure becomes negative, the system (2.28) will be non-hyperbolic and thus the initial value problem will be ill-posed. Here, we try to utilize the positivity-preserving (PP) limiter [41] to avoid this problem. However, our time discretization is different from the standard SSP-RK method, which is used in [41]. Hence, we need to reanalysis the limiter to match the IF-SSP-RK scheme.
Firstly, we give a brief review of the PP limiter [41] for one-dimensional Euler equation (2.28) based on the standard DG scheme coupling with the forward Euler time discretization
$\int_{K_{i}}{\frac{\mathbf{u}_{h}^{n+1}-\mathbf{u}_{h}^{n}}{\Delta t}\cdot\mathbf{w}\mathrm{d}x}-\int_{K_{i}}{\mathbf{f}\left(\mathbf{u}_{h}^{n}\right)\cdot\frac{\partial\mathbf{w}}{\partial x}\mathrm{d}x}+\mathbf{\hat{f}}_{i+1/2}^{n}\mathbf{w}_{i+1/2}^{-}-\mathbf{\hat{f}}_{i-1/2}^{n}\mathbf{w}_{i-1/2}^{+}=\mathbf{0}.$ (3.40)
Let $\mathcal{G}\subset\mathbb{R}^{3}$ denote the admissible set
$\mathcal{G}:=\left\{(\rho,m,\mathcal{E}):\rho>0\ \ \mathrm{and}\ \ p=(\gamma-1)\left(\mathcal{E}-\frac{1}{2}\frac{m^{2}}{\rho}\right)>0\right\}$
which is convex. It was proved that the first order DG scheme ($k=0$) is positivity preserving under a certain CFL condition $\frac{\Delta t}{h}\left\|(\left|u\right|+c)\right\|_{L^{\infty}(\Omega)}\leq\alpha_{0}$, that is if $\bar{\mathbf{u}}_{i}^{n}\in\mathcal{G}$, then $\bar{\mathbf{u}}_{i}^{n+1}\in\mathcal{G}$. Here, $\bar{\mathbf{u}}_{i}^{n}$ is the cell average of the numerical solution in the cell $K_{i}$ at time $t_{n}$.
[…]
Moreover, [41] showed that the standard high order DG scheme can maintain $\bar{\mathbf{u}}_{i}^{n+1}\in\mathcal{G}$ under a suitable CFL condition, as long as the values of the approximation polynomials at the Gauss-Lobatto quadrature points belong to $\mathcal{G}$. Based on these conclusions, authors designed a PP limiter as a simplified linear scaling limiter, without destroying the original order of accuracy. Since a SSP-RK method is a convex combination of Euler steps, the positivity-preserving can still hold if we apply the limiter on each stage of SSP-RK methods.
In the following, we will design the PP-limiter on the NOES-DG scheme (2.31). Let $\{(\hat{x}_{i}^{q},\omega_{q})\}_{q=1}^{k+1}$ be the Gauss-Lobatto points and quadrature weights on a interval $K_{i}$. Note that the NOES-DG scheme coupling with Euler forward can generate the following form
@@
which is the same as that of the traditional DG scheme. This tell us that the artificial viscosity term does not affect the cell-averages. Hence, we can obtain the same conclusion for the NOES-DG scheme.
Theorem 3.1.
Consider the NOES-DG scheme coupling with Euler forward discretization. If the numerical solution $\mathbf{u}_{h}^{n}$ satisfies
$\mathbf{u}_{h}^{n}(\hat{x}_{i}^{q})\in\mathcal{G},\quad\forall i,q.$ (3.41)
Then $\bar{\mathbf{u}}_{i}^{n+1}\in\mathcal{G}$ under the CFL condition
$\frac{\Delta t}{h}\left\|(\left|u\right|+c)\right\|_{L^{\infty}(\Omega)}\leq\omega_{1}\alpha_{0}.$ (3.42)
The proof is the same as that of the traditional DG scheme [41], which would not been repeated here.
Furthermore, based on the consistency of these two methods, we can utilized the the PP-limiter in [41] on our NOES-DG scheme to meet the condition (3.41), which is realized with the help of two scaling parameters $\theta_{i}^{(1)},\theta_{i}^{(2)}\in[0,1]$ on each cell $K_{i}$, and set
@@
to guarantee $\tilde{\rho}_{h}^{n}(\hat{x}^{q}_{i})\geq 0$ and $\mathbf{u}_{h}^{n,(mod)}(\hat{x}^{q}_{i})\in\mathcal{G}$. To be specific, the value of $\theta$ is calculated by
@@
where $\tilde{\mathbf{u}}_{h}=(\tilde{\rho}_{h}^{n},m_{h}^{n},\mathcal{E}_{h}^{n})$, and $\tilde{t}_{q}$ satisfies
$p((1-\tilde{t}_{q})\bar{\mathbf{u}}_{i}+\tilde{t}_{q}\tilde{\mathbf{u}}_{h}(\hat{x}_{i}^{q}))=\varepsilon$
can be obtained by solving a nonlinear equation. The parameter $\varepsilon$ is a small number, and we set $\varepsilon=10^{-13}$ as the same in [41]. We want to remark that the scaling PP-limiter (3.43) does not change the cell average, and moreover, as showed in [41], it can maintain the original high order of accuracy for smooth solutions.
Next, we concern on the proposed NOES-DG scheme coupling with the IF-SSP-RK scheme, and have the following conclusion.
Theorem 3.2.
Consider the NOES-DG scheme (2.31) with PP limiter (3.43) and IF-SSP-RK time discretization (3.37), then the numerical solution satisfies $\bar{\mathbf{u}}_{i}^{n+1}\in\mathcal{G}$ under the CFL condition
$\frac{\Delta t}{h}\left\|(\left|u\right|+c)\right\|_{L^{\infty}(\Omega)}\leq\frac{3}{4}\omega_{1}\alpha_{0}.$ (3.46)
Proof.
We denote the PP limiter as an operator $\mathcal{M}$. Thus, the fully discrete scheme yields
$\mathbf{\bar{u}}_{i}^{\left(1\right)}$ $=\frac{1}{2}\mathbf{\bar{u}}_{i}^{n}+\frac{1}{2}\left(\mathbf{\bar{u}}_{i}^{n}+\frac{4}{3}\Delta t\cdot\mathcal{L}_{i}^{ave}\left(\mathcal{M}\left(\mathbf{u}_{h}^{n}\right)\right)\right),$ (3.47)
$\mathbf{\bar{u}}_{i}^{\left(2\right)}$ $=\frac{2}{3}\mathbf{\bar{u}}_{i}^{n}+\frac{1}{3}\left(\mathbf{\bar{u}}_{i}^{\left(1\right)}+\frac{4}{3}\Delta t\cdot\mathcal{L}_{i}^{ave}\left(\mathcal{M}\left(\mathbf{u}_{h}^{\left(1\right)}\right)\right)\right),$
$\mathbf{\bar{u}}_{i}^{n+1}$ $=\frac{59}{129}\mathbf{\bar{u}}_{i}^{n}+\frac{15}{129}\left(\mathbf{\bar{u}}_{i}^{n}+\frac{4}{3}\Delta t\cdot\mathcal{L}_{i}^{ave}\left(\mathcal{M}\left(\mathbf{u}_{h}^{n}\right)\right)\right)+\frac{27}{64}\left(\mathbf{\bar{u}}_{i}^{\left(2\right)}+\frac{4}{3}\Delta t\cdot\mathcal{L}_{i}^{ave}\left(\mathcal{M}\left(\mathbf{u}_{h}^{\left(2\right)}\right)\right)\right).$
It can be seen that (3.47) is also a convex combination of Euler steps, but the time step is changed to $34\Delta t/$ 3. Since the PP limiter ensures $\mathcal{M}(\mathbf{u}^{n}_{h})(\hat{x}_{i}^{q})\in\mathcal{G}$ for each $q$, thus under the CFL condition (3.46), we have $\bar{\mathbf{u}}_{i}^{(1)}\in\mathcal{G}$. Likewise, we can further have $\bar{\mathbf{u}}_{i}^{(2)}\in\mathcal{G}$ and $\bar{\mathbf{u}}_{i}^{n+1}\in\mathcal{G}$. ∎
@@
For the two-dimensional cases, the PP collocation of a rectangular cell is changed to $\mathbb{S}_{2D}=(\mathbb{S}_{i}^{x}\times\hat{\mathbb{S}}_{j}^{y})\bigcup(\hat{\mathbb{S}}_{i}^{x}\times\mathbb{S}_{j}^{y})$, where $\mathbb{S}_{i}^{x}$ and $\hat{\mathbb{S}}_{i}^{x}$ denote the Gauss points and Gauss-Lobatto points on $[x_{i-1/2},x_{i+1/2}]$, respectively. Then the scaling coefficients $\theta^{(1)}_{ij},\theta^{(2)}_{ij}$ are calculated on these $2(k+1)^{2}$ points [41]. For triangular meshes, the main difference is of the scaling limiter points collocation, and the details can refer to [42].
@@ 4 Numerical tests
and $\mathrm{CFL}$ is taken as $0.2$. Without special declaration, we take the parameter $C=50$ for all tests. And $c_{f}$ in $\sigma_{i}^{jump}$ is defined as $c_{f}=c_{0}\max\limits_{\mathbf{x}\in K}\{\left|\nabla\cdot\mathbf{f}\right|\}$ and $c_{f}=c_{0}\max\limits_{\mathbf{x}\in K}\{1/H\}$ in cell $K$ for scalar equations and Euler system respectively, where $H=(\mathcal{E}+p)/\rho$ denotes the enthalpy. Time evolution of the total entropy will be given for some problems, which is defined as $\int_{\Omega}U(u_{h})\mathrm{d}x$.
@@ 4.1 One-dimensional tests
First, we test the accuracy of the scheme with the smooth initial data $u(x,0)=\sin(x)$. In Table 4.1, we present the errors and orders of accuracy at $T=2\pi$. We can observe the optimal convergence rates for $k=1,2,3$.
@@ 4.3 Euler system
In Figure 4.9, we compare the result of density at $T=0.2$ with exact solution on $N=200$ meshes. We can see the results are well-resolved and nonoscillatory.
@@
Here, we consider the two blast-waves problem [40], which is very challenging to most of the high-order schemes. The computational domain is $\Omega=[0,1]$ with reflective boundary conditions imposed on both end points, and the initial data is given by
@@
In Figure 4.11, we present the result of density at $T=0.038$ with $N=800$ cells. The reference solution is obtained with the fifth-order WENO scheme with 2000 cells. Notice that the classical DG scheme suffers from negative pressure or negative density in this examples. However, our proposed scheme with PP limiter performs well.
@@
Here, we consider the Leblanc problem [31], which is an extremely Riemann problem with very strong discontinuities and is used to test robustness of our scheme. The computational domain is $\Omega=[-10,10]$, and the initial data is given by
@@
In Figure 4.12, we compare the result of density, velocity and pressure with the exact solution at $T=0.0001$ on $N=800$ and $N=6400$ meshes. These results show the robustness and convergence of our scheme.
@@
$\left(\rho,u,v,p\right)=\begin{cases}\left(8,8.25\cos\left(\frac{\pi}{6}\right),8.25\sin\left(\frac{\pi}{6}\right),116.5\right),&x<\frac{1}{6}+\frac{y}{\sqrt{3}},\\ \left(1.4,0,0,1\right),&x\geq\frac{1}{6}+\frac{y}{\sqrt{3}}.\\ \end{cases}$
@@
A Mach 2000 jet state $(\rho,u,v,p)=(5,800,0,0.4127)$ is injected from the left boundary in the range $-0.05<y<0.05$. Outflow conditions are applied on all remaining boundaries.
@@ 5 Concluding remarks
In this paper, we present a class of non-oscillatory entropy-stable DG methods for solving hyperbolic conservation laws. By incorporating a specific form of local artificial viscosity term into the original DG scheme on each cell and imposing an interpolation condition, the semi-discrete scheme ensures strict non-increase of the total entropy with a careful design of the viscosity coefficient. An oscillation coefficient is taken into account to control spurious oscillation near shock waves. Moreover, the oscillation coefficient is also used to limit viscosity parameter for cases where the denominator of the entropy coefficient approaches zero. We prove the optimal accuracy for scalar advection equations with quadratic entropy, and advection equations with other entropy functions under certain assumptions. It is noted that large viscosity terms can induce stiffness and impose strict constraints on time step sizes. To address this, we employ the integration factor Runge-Kutta method, enabling explicit computation under standard CFL conditions. Due to the local structure of our proposed scheme, the algorithm can be implemented easily. For extreme problems where negative pressures might occur, we prove that our scheme is compatible with positivity-preserving limiters under suitable CFL conditions. Extensive numerical examples confirm that our scheme guarantees entropy decrease, controls oscillations, and achieves optimal accuracy. Future work may include extending this method to unstructured meshes and applying it to other equations.

### v2

@@ 3.1 Integral factor SSP-RK method
The PDE can be solved in the MOL framework, that is after DG discretization the semi-discrete scheme can be treated as an ODE system, and then, we can apply a known ODE solver on it to obtain the fully discrete scheme. However, for numerical solutions containing discontinuities, the additional viscous term becomes stiff, resulting in a severely restricted time step when using explicit time discretization. To overcome this difficulty, we will employ the integrating factor strong-stability-preserving Runge-Kutta (IF-SSP-RK) method introduced in [6]. Here, we only discuss the one-dimensional scalar case for brevity.
@@
To introduce the IF-SSP-RK method, we begin with a specific case that this stiff term is semilinear, i.e. $\mathcal{L}^{v}_{h}(\mathbf{u}_{h})=\mathbf{C}\mathbf{u}_{h}$, where $\mathbf{C}$ is a constant matrix. Then we can multiply the factor $e^{-\mathbf{C}t}$ on both sides and get that
@@ 3.2 Positivity-preserving limiter
For certain extreme cases, negative density or pressure may appear in the numerical solution produced by the proposed method. In such cases, the system (2.28) becomes non-hyperbolic, rendering the initial value problem ill-posed. To overcome the difficulty, we propose to utilize the PP limiter [42]. Denote $\bar{\mathbf{u}}_{i}^{n}$ as the cell average of the numerical solution $u_{h}$ in the cell $K_{i}$ at time $t_{n}$:
$\bar{\mathbf{u}}_{i}^{n}=\frac{1}{h}\int_{K_{i}}\mathbf{u}_{h}(x,t^{n})\mathrm{d}x.$
Note that the cell averages satisfy the same equation as in the standard RKDG formulation since the elements in the corresponding positions of $\mathbf{C}$ are all zero. Therefore, although the time discretization used in this work differs from the standard SSP-RK method employed in the original work [42], both approaches are essentially equivalent when applying the PP limiter.
First, we briefly review the PP limiter developed in [42] for the one-dimensional Euler equations (2.28) based on the standard SSP-RKDG method. Let $\mathcal{G}\subset\mathbb{R}^{3}$ denote the admissible set
$\mathcal{G}:=\left\{(\rho,m,\mathcal{E}):\rho>0\ \ \mathrm{and}\ \ p=(\gamma-1)\left(\mathcal{E}-\frac{1}{2}\frac{m^{2}}{\rho}\right)>0\right\},$
which is convex. It was proved that the first order DG scheme ($k=0$) coupled with the forward Euler method is PP under a certain CFL condition $\frac{\Delta t}{h}\left\|(\left|u\right|+c)\right\|_{L^{\infty}(\Omega)}\leq\alpha_{0}$. That is if $\bar{\mathbf{u}}_{i}^{n}\in\mathcal{G}$, then $\bar{\mathbf{u}}_{i}^{n+1}\in\mathcal{G}$. Moreover, [42] shows that a high order DG scheme, when coupled with the forward Euler method, can ensure $\bar{\mathbf{u}}_{i}^{n+1}\in\mathcal{G}$ under a suitable CFL condition, provided that the nodal values of the approximate polynomials at the Gauss-Lobatto quadrature points remain within $\mathcal{G}$. Building on this, authors designed a simple scaling-based PP limiter that preserves the original order of accuracy. Since an SSP-RK method is a convex combination of forward Euler steps, the PP property still holds for an SSP-RKDG method as long as the limiter is applied at each stage of the SSP-RK method.
In the following, we will apply the PP-limiter on the NOES-DG scheme (2.31). Let $\{(\hat{x}_{i}^{q},\hat{\omega}_{i}^{q})\}_{q=1}^{k+1}$ be the Gauss-Lobatto points and quadrature weights on a interval $K_{i}$. Note that the cell average of the numerical solution by the NOES-DG scheme coupled with the IF forward Euler method satisfies the following formulation
@@
which is the same as that of the traditional DG scheme coupled with the forward Euler method. Hence, based on such equivalence, we can directly employ the PP-limiter in [42] for our NOES-DG scheme. In particular, the DG solution is rescaled using two scaling parameters $\theta_{i}^{(1)},\theta_{i}^{(2)}\in[0,1]$ on each cell $K_{i}$:
@@
to guarantee $\tilde{\rho}_{h}^{n}(\hat{x}^{q}_{i})\geq 0$ and $\mathbf{u}_{h}^{n,(mod)}(\hat{x}^{q}_{i})\in\mathcal{G}$. The values of $\theta_{i}^{(1)}$ and $\theta_{i}^{(2)}$ are given by
@@
where $\tilde{\mathbf{u}}_{h}=(\tilde{\rho}_{h}^{n},m_{h}^{n},\mathcal{E}_{h}^{n})$, and $\tilde{t}_{q}$ is solved from the following nonlinear equation
$p((1-\tilde{t}_{q})\bar{\mathbf{u}}_{i}+\tilde{t}_{q}\tilde{\mathbf{u}}_{h}(\hat{x}_{i}^{q}))=\varepsilon,$
where the parameter $\varepsilon$ is a small number, and we set $\varepsilon=10^{-13}$ as suggested in [42]. Finally, we note that the scaling PP-limiter (3.42) maintains the cell averages. Moreover, as shown in [42], it preserves the original high order of accuracy for smooth solutions.
@@
For the two-dimensional case, the quadrature points for the PP limiter on a rectangular cell are defined as $\mathbb{S}_{2D}=(\mathbb{S}_{i}^{x}\times\hat{\mathbb{S}}_{j}^{y})\bigcup(\hat{\mathbb{S}}_{i}^{x}\times\mathbb{S}_{j}^{y})$, where $\mathbb{S}_{i}^{x}$ and $\hat{\mathbb{S}}_{i}^{x}$ denote the Gauss quadrature points and Gauss-Lobatto quadrature points on $[x_{i-1/2},x_{i+1/2}]$, respectively. The scaling parameters $\theta^{(1)}_{ij},\theta^{(2)}_{ij}$ are calculated at these $2(k+1)^{2}$ points, see [42]. For triangular meshes, the main difference lies in the quadrature rule, and the details can be found in [43].
@@ 4 Numerical tests
and $\mathrm{CFL}$ is taken as $0.2$. Unless otherwise noted, we take the parameter $C=50$ for all tests. Moreover, $c_{0}$ is a free parameter, and the optimal value is problem-dependent. Our selection principle is to choose a small value while preventing spurious oscillations. Based on our experience, a suitable range is $100.1<c_{0}<$ 10. Further investigation into these parameters, including the scale-invariant properties, will be considered in future work. Time evolution of the total entropy will be given for some problems, which is defined as $\int_{\Omega}U(u_{h})\mathrm{d}x$.
@@ 4.1 One-dimensional tests
First, we test the accuracy of the scheme with the smooth initial data $u(x,0)=\sin(x)$. In Table 4.1, we present the errors and orders of accuracy at $T=2\pi$. We can observe the optimal convergence rates for $k=1,2,3$. For this test problem, we also provide the CPU cost comparison between the proposed NOES-DG scheme and standard DG scheme for $k=2$ in Table 4.2. It is observed that the percentage of increase in cost is negligible (less than 4%), highlighting the efficiency of the proposed technique.
@@
Table 4.2: Example 4.1: One dimensional linear equation with a smooth initial data. The comparison of CPU cost (in seconds) for $k=2$ up to $T=2\pi$.
$N$ Standard NOES Ratio
64 0.3996s 0.4116s 1:1.0300
128 1.3884s 1.4434s 1:1.0396
256 5.3796s 5.4651s 1:1.0159
512 21.1410s 21.2178s 1:1.0036
@@ 4.3 Euler system
[…]
@@
(a) Density with different $c_{0}$.
(b) Zoomed in results.
Figure 4.10: Example 4.7: Sod shock tube. The numerical solution at $T=0.2$ with $N=200$ with different values of $c_{0}$.
@@
Here, we consider the two blast-waves problem [41], which presents a significant challenge for most high-order schemes. The computational domain is $\Omega=[0,1]$ with reflective boundary conditions imposed on both end points, and the initial data is given by
@@
The PP limiter is needed to ensure positivity of pressure and density. In Figure 4.12, we present the result of density at $T=0.038$ with $N=800$ cells. The reference solution is obtained with the fifth-order WENO scheme with 2000 cells. The proposed scheme produces high-quality numerical results with sharp, non-oscillatory shock transitions.
@@
Here, we consider the Leblanc problem [31], which is an extreme Riemann problem with very strong discontinuities and is used to test robustness of our scheme. The computational domain is $\Omega=[-10,10]$, and the initial data is given by
@@
This example, like the previous two blast waves test, requires the PP limiter. In Figure 4.13, we compare the result of density, velocity and pressure with the exact solution at $T=0.0001$ on $N=800$ and $N=6400$ meshes. These results show the robustness and convergence of our scheme.
@@
$\left(\rho,u,v,p\right)=\begin{cases}\left(8,8.25\cos\left(\frac{\pi}{6}\right),-8.25\sin\left(\frac{\pi}{6}\right),116.5\right),&x<\frac{1}{6}+\frac{y}{\sqrt{3}},\\ \left(1.4,0,0,1\right),&x\geq\frac{1}{6}+\frac{y}{\sqrt{3}}.\\ \end{cases}$
@@
A Mach 2000 jet state $(\rho,u,v,p)=(5,800,0,0.4127)$ is injected from the left boundary in the range $-0.05<y<0.05$. Outflow conditions are applied on all remaining boundaries. This example requires the PP limiter.
@@ 5 Concluding remarks
In this paper, we present a class of non-oscillatory entropy-stable DG methods for solving hyperbolic conservation laws. By incorporating a specific form of local artificial viscosity term into the original DG scheme on each cell and imposing an interpolation condition, the semi-discrete scheme ensures strict non-increase of the total entropy with a careful design of the viscosity coefficient. A damping coefficient is taken into account to control spurious oscillation near shock waves. Moreover, this damping coefficient is also employed to constrain the viscosity parameter in cases where the entropy coefficient’s denominator approaches zero. We prove the optimal accuracy for scalar advection equations with quadratic entropy, and advection equations with other entropy functions under certain assumptions. It is noted that the added viscosity term may induce stiffness and impose strict constraints on time step sizes. To address this, we employ the integrating factor Runge-Kutta method, enabling explicit computation under the standard CFL condition. Due to the local structure of our proposed scheme, the algorithm can be implemented easily. For extreme problems where negative pressures might occur, we demonstrate that our scheme is compatible with the positivity-preserving limiter. Extensive numerical examples confirm that our scheme ensures entropy stability, suppresses oscillations, and attains optimal accuracy. Future work involves extending the proposed method to unstructured meshes and exploring its application to other equations.